The mean caloric intake of males aged 20-39 years old is 2716 calories, with a standard deviation of 72.8 calories. Suppose a nutritionist analyzes a simple random sample of 35 males aged 20-39 and finds that their mean intake was 2750 calories. What is the probability that a random sample of 35 males aged 20-39 would result in a sample mean of 2750 calories or higher? Show All Work Pls.
Answer #1

Solution :

Given that ,

mean = \mu = 2716

standard deviation = \sigma = 72.8

n = 35

\mu\bar x = 2716

\sigma\bar x = \sigma / \sqrt n = 72.8 / \sqrt35 = 12.305

P(\bar x >2750 ) = 1 - P(\bar x < 2750)

= 1 - P[(\bar x - \mu \bar x ) / \sigma \bar x < (2750-2716) /12.305 ]

= 1 - P(z <2.76 )

Using z table

= 1 - 0.9971

= 0.0029

probability= 0.0029